4. Kaitlin Block Publishing, a textbook publishing firm, purchased a new machine for $120,000. This machine is expected to operate for 10 years, after which it will be sold for salvage value (estimated to be $9,000). How much will the first and second year’s depreciation expense be under the double-declining-balance method?

5. Scheuller Company had machinery that had originally cost $246,000. The machinery was three years old and had been depreciated using the double-declining-balance method, over a five-year useful life with a residual value of $18,000. Answer each of the following independent questions: Required: a. If the company sold the machinery for $105,000, prepare a journal entry to record the sale. b. If the company sold the machinery for $48,000, prepare a journal entry to record the sale.
Answer #1

4. Double declining rate = 200/useful life = 200/10 = 20% per year

Depreciation for

First year = 120,000 * 20% = 24,000

Second Year = (120,000-24,000) * 20% = 19,200

As per HomeworkLib guidelines I'm supposed to solve your first question.